### 15 Apr 2026 v3

# Investing is Compression

Abstract.
In 1956 John Kelly wrote a paper at Bell Labs describing the relationship between gambling and Information Theory. What came to be known as the Kelly Criterion is both an objective and a closed-form solution to sizing wagers when odds and edge are known. Samuelson argued it was arbitrary and subjective, and successfully kept it out of mainstream economics. Luckily it lived on in computer science, mostly because of Tom Cover’s work at Stanford. He showed that it is the uniquely optimal way to invest: it maximizes long-term wealth, minimizes the risk of ruin, and is competitively optimal in a game-theoretic sense, even over the short term.
One of Cover’s most surprising contributions to portfolio theory was the universal portfolio. Related to universal compression in information theory, it performs asymptotically as well as the best constant-rebalanced portfolio in hindsight. I borrow a trick from that algorithm to show that Kelly’s objective, even in the general form, factors the investing problem into three terms: a money term, an entropy term, and a divergence term. The only way to maximize growth is to minimize divergence which measures the difference between our distribution and the true distribution in bits. Investing is, fundamentally, a compression problem.
This decomposition also yields new practical results. Because the money and entropy terms are constant across strategies in a given backtest, the difference in log growth between two strategies measures their relative divergence in bits. I also introduce a winner fraction heuristic which allocates capital in proportion to each asset’s probability of dominating the candidate set. The growth shortfall of this heuristic relative to the optimal portfolio is bounded by the entropy of the winner fraction distribution. To my knowledge, both the heuristic and the entropy bound are original contributions.
1. Introduction
In 1956 John Kelly wrote a paper while working at Bell Labs [9] describing the relationship between gambling and Information Theory. The paper describes what would eventually be known as the Kelly criterion. The term Kelly criterion has come to signify two distinct things: the objective or utility function, as well as the closed-form solution to the problem when a gambler’s odds and edge are known. Kelly observed that simply maximizing expected wealth would surely lead to ruin given enough time. What was needed was some kind of nonlinear utility function to balance risk and reward. The math of utility functions came from Von Neumann and Morgenstern’s game theory. It formalizes a subjective preference or willingness to risk loss in pursuit of gain. Kelly wanted to avoid the “introduction of an arbitrary cost function,” likening it to the external value that people attempted to attach to Shannon’s information. Kelly realized that maximizing logarithmic growth rather than simple expectation simultaneously avoids ruin while maximizing asymptotic wealth. Kelly wasn’t the first to propose this objective. It was originally proposed by Bernoulli in 1738 as the solution to the Saint Petersburg paradox[4]. But Kelly was the first to connect it to the new field of information theory, encouraged by Shannon himself who assisted with the paper and immediately realized the significance.
It wasn’t clear at the time what was so special about this particular function. It looks a little bit like Shannon’s celebrated entropy equation, but with money replacing the term inside the logarithm. The function balances potential gains and losses along with their probabilities through the curvature of the log function. Because log is steeper to the left than to the right at every point, losses matter more than gains. And log of zero is negative infinity, so it’s not enough to have a small probability of blowup (ruin) – it has to be strictly zero. Another interpretation is that investments should be both structurally and statistically asymmetric.
Shannon shared the paper with Edward Thorp, who applied it first to blackjack [18], then to the stock market through Princeton-Newport Partners, building one of the earliest quantitative investment funds and achieving one of the most successful track records in investing history [19]. Thorp also conducted extensive mathematical analysis of the Kelly criterion, extending it to practical multi-asset and options settings. [11]
Shannon separately presented these ideas to the economist Paul Samuelson at MIT. Over the course of several evening meetings [7], they discussed and debated the merits of this proposal. Samuelson is said to have had a “visceral dislike of log utility.” He argued it was an arbitrary and subjective utility function, and an unnecessarily conservative one at that. Over the course of several papers, he defended economics against this incursion by computer science and “fallacy that has been borrowed into portfolio theory from information theory of the Shannon type.”[14, 13, 15] As perhaps the most influential economist at the time, he successfully kept this idea out of mainstream economics.
Thankfully, this idea lived on in computer science, most notably because of the work of Tom Cover, the Stanford information theory professor. We now know because of Cover’s work that Shannon’s idea was not a subjective, arbitrary utility function as Samuelson had argued. It is the objectively optimal way to invest for three reasons.
  1. (1)
    It maximizes long term wealth.
  2. (2)
    It minimizes the risk of ruin.
  3. (3)
    In a game-theoretic sense, it is competitively optimal, even over the short term. Not just asymptotically as Samuelson believed. [2, 3, 1]
Samuelson seems to have understood the first two points, but he remained unconvinced, believing there was nothing special about log that distinguished it from other utility functions. The third reason, the most important, wasn’t known until much later because of Cover’s work at Stanford. Proving optimality required new ideas and techniques in information theory, including a version of the Asymptotic Equipartition Property (AEP) for markets to show that it dominates every other strategy over the long run [1], and an application of Kuhn-Tucker conditions [3] to show that it is also optimal in a game-theoretic sense over the short term. Cover noted that it was “tantalizing that [the log optimal strategy] arises as the solution to such dissimilar problems” and wondered if it was just a mathematical coincidence or something more fundamental. [2]. As Cover observed:
“there is no essential conflict between good short-term and long-run performance. Both are achieved by maximizing the conditional expected log return.” Cover [3]
One of Cover’s most surprising contributions to portfolio theory was something called a universal portfolio. Related to universal compression in information theory, it performs asymptotically as well as the best possible adversary. One that knows the rules but not the future. Although the math is elegant, the argument is very abstract. But one of the steps that Cover used in creating universal portfolios is extremely useful beyond that algorithm. It reveals the information structure of the investing problem, making it accessible to the techniques of information theory.
That same technique will be applied here to show that what is so powerful about the Kelly criterion is that even in the most general form, it factors the investing problem into three terms: a money term, an entropy term, and a divergence term. The first two terms don’t actually depend on our allocation or distribution. So the only way to maximize our compounding growth rate is to minimize the divergence which is a non-negative drag or friction and which only goes to zero when our distribution matches the unknown true distribution. This means that, fundamentally, investing is a compression problem.
2. Classic Kelly
The classic Kelly criterion tells us how much to wager on a game of chance with known odds and predictive edge. Assume that with probability $p$ we win and our wager is multiplied by return $r$. And with probability (1p)(1-p) we lose the full wager. The question is what fraction $f$ to wager to maximize wealth. Note that $r$ is the return multiplier, not the more common $\text{odds}=r-1$.
If we start with initial wealth $V_{0}$, then with probability $p$ we will win and our wealth at the next step will be
(1) $V_{1}=V_{0}(fr+(1-f))$
and with probability (1p)(1-p) we will lose and our wealth will be
(2) $V_{1}=V_{0}(1-f)$
After n steps, our wealth becomes
(3) $V_{n}=V_{0}(1-f)^{n_{0}}(fr+(1-f))^{n_{1}}$
if we have $n_{0}$ losses and $n_{1}$ wins.
Kelly’s proposal is that we maximize the expected log wealth, or the probability weighted average log return. Note that log is monotonic, so maximizing log return also maximizes total growth after $n$ steps. Kelly’s focus was the expected log-growth per step, $g$. Since each step is multiplicative, the log of the full product is the sum of the logs, and the average converges to the expectation in the limit.
(4) $g(f)=\lim_{n\to\infty}\frac{\log\left(\frac{V_{n}}{V_{0}}\right)}{n}=p\log(fr+(1-f))+(1-p)\log(1-f)$
We can find the $f^{*}$ that maximizes g by taking the derivative and setting it equal to zero
(5) $\frac{dg}{df}=\frac{p(r-1)}{fr-f+1}-\frac{1-p}{1-f}=0$
Solving for f we get
(6) $f^{*}=\frac{pr-1}{r-1}$
This is the famous “edge/odds” Kelly solution that maximizes growth. Plugging it back into the growth equation:
(7) $g^{*}=p\log(pr)+(1-p)\log\left(\frac{(1-p)r}{r-1}\right)$
If the market maker’s probability is $q$ and the odds are fair, then $q\cdot r=1$, and $r=\frac{1}{q}$. Substituting into the equation:
(8) $g^{*}=p\log\left(\frac{p}{q}\right)+(1-p)\log\left(\frac{1-p}{1-q}\right)$
This is the KL divergence $\operatorname{D_{KL}}(P\|Q)$ between our distribution p and the market’s distribution q.
3. Horse Race Market
Consider a market where you can bet fractionally on mutually exclusive events and you must fully allocate your capital. This is described as a “horse race market” [5] with only one winner each time. In this case we need a weight vector $W$ rather than a single wager fraction $f$. If there are $n$ horses, let $w_{i}$ be the fraction wagered on the $i$th horse, let $p_{i}$ be the probability of the $i$th horse winning, and let $r_{i}$ be the return or the wealth ratio for capital wagered on that horse. Kelly’s expected log return for this game becomes:
(9) $g(W)=\sum_{i=1}^{N}p_{i}\log\left(r_{i}w_{i}\right)$
In the classic Kelly example, we used calculus to find the optimal solution and observed the information theoretic KL divergence in the solution. This time we will take a more direct information theoretic approach to find and characterize the solution. This problem is simple to analyze because there is only one nonzero term inside the logarithm (the single winning horse).
Because the log of a product is the sum of logs, we can rewrite this as
(10) $g(W)$ $=\sum_{i=1}^{N}p_{i}\log(r_{i}w_{i})$
(11) $=\sum_{i=1}^{N}p_{i}\log(r_{i})+\sum_{i=1}^{N}p_{i}\log\left(w_{i}\right)$
(12) $=\sum_{i=1}^{N}p_{i}\log(r_{i})+\sum_{i=1}^{N}p_{i}\log\left(\frac{p_{i}w_{i}}{p_{i}}\right)$
(13) $=\sum_{i=1}^{N}p_{i}\log(r_{i})+\sum_{i=1}^{N}p_{i}\log\left(p_{i}\right)+\sum_{i=1}^{N}p_{i}\log\left(\frac{w_{i}}{p_{i}}\right)$
(14) $=\sum_{i=1}^{N}p_{i}\log(r_{i})-\sum_{i=1}^{N}p_{i}\log\left(\frac{1}{p_{i}}\right)-\sum_{i=1}^{N}p_{i}\log\left(\frac{p_{i}}{w_{i}}\right)$
(15) $=\sum_{i=1}^{N}p_{i}\log\left(r_{i}\right)-\operatorname{H}(P)-\operatorname{D_{KL}}(P\|W)$
Notice that the first two terms don’t depend on W. The only way to maximize growth is to minimize the third term which is the KL divergence between the probability distribution and the weight vector. Because the weight vector sums to one, we can think of it like a probability distribution. We know from Gibbs’ Inequality that $\operatorname{D_{KL}}(P\|W)$ is non-negative and only goes to zero when $P=W$. In other words, the optimal strategy is to bet in proportion to the true probability distribution. This is a counterintuitive fact because it does not depend on the return distribution, just the probability of winning.
If the payoffs are priced fairly using a market probability Q, then as before $q_{i}\cdot r_{i}=1$ and so $q_{i}=\frac{1}{r_{i}}$.
Substituting that back into the above solution, we have
(16) $g(W)$ $=\sum_{i=1}^{N}p_{i}\log\left(\frac{1}{q_{i}}\right)-\operatorname{H}(P)-\operatorname{D_{KL}}(P\|W)$
(17) $=\operatorname{H}(P,Q)-\operatorname{H}(P)-\operatorname{D_{KL}}(P\|W)$
(18) $=\operatorname{D_{KL}}(P\|Q)-\operatorname{D_{KL}}(P\|W)$
Which follows from the fact that KL divergence is the difference between the cross entropy and entropy.
The growth rate is the difference between these two KL divergences: the divergence of the market from the true distribution minus the divergence of our allocation from the truth. The intuition is that you can’t make money if the market is priced accurately or if you simply agree with the market. You have to disagree with the market while being closer to the true distribution.
4. General Case
The Horse Race Market entails mutually exclusive events, with only one nonzero term inside the logarithm at each step. This makes it straightforward to analyze using standard mathematical techniques. Consider now the more general case where there can be more than one nonzero term at each point in time. In this case, the return at each point in time is the inner product between the weight vector and the return vector:
(19) $\sum_{i=1}^{N}r_{i}w_{i}$
and the log return is given by
(20) $\log\left(\sum_{i=1}^{N}r_{i}w_{i}\right)$
The Horse Race Market expected log return
(21) $\sum_{i=1}^{N}p_{i}\log(r_{i}w_{i})$
is a special case of this more general formulation in which the ith term is nonzero with probability $p_{i}$. The general case is more difficult to analyze because the sum doesn’t have the same clean factorization.
To analyze this problem, I’m going to use a technique that was first used by Tom Cover in his universal portfolio theory [6, 5] proofs. Although Cover used the technique for a different purpose, it turns out to be extremely useful for analyzing the general investing problem. It’s essentially a change of representation that reveals an information structure to the investing problem and makes it more accessible to techniques and insights from information theory. Cover used this trick to make it possible to apply the math of universal compression which is the surprising idea from information theory that you can achieve the Shannon compression limit asymptotically even without knowing the probability distribution ahead of time. In the context of investing, this seems especially paradoxical, but the math is nevertheless sound and elegant. Rather than using it to design an algorithm as Cover did, I’m simply borrowing the technique for the purpose of analysis.
We traditionally think of the multi-period investing problem as a product of scalars. We can write the return after $n$ periods of some multi-period investment process as the product of the individual returns:
(22) $R_{n}=\prod_{t=1}^{n}r_{t}$
where $R_{n}$ indicates the total growth of the multiplicative process after $n$ steps.
Assuming we are compounding and rebalancing a fixed weight portfolio at each point in time, we can write this full process, where $w_{i}$ is the portfolio weight assigned to the ith asset, and $r_{i,t}$ is the return of the ith asset at time $t$.
(23) $R_{n}=\prod_{t=1}^{n}\sum_{i=1}^{m}w_{i}r_{i,t}$
Note that the product of scalars is now seen as a product of sums (dot products).
Cover’s brilliant insight was that we can rewrite the product of sums as a sum of products. This might seem like a trivially algebraic change. But as will become clear momentarily, instead of compounding weighted returns per period, we can think of this as a weighted average of an exponential number of multi-period returns.
Assume for now that there are only two assets in the system $\{a,b\}$. The return after one period is
(24) $w_{a}r_{a,1}+w_{b}r_{b,1}$
After two periods the return is
(25) $\left(w_{a}r_{a,1}+w_{b}r_{b,1}\right)\left(w_{a}r_{a,2}+w_{b}r_{b,2}\right)$
Which we can expand as
(26) $w_{a}^{2}r_{a,1}r_{a,2}+w_{b}^{2}r_{b,1}r_{b,2}+w_{a}w_{b}r_{a,1}r_{b,2}+w_{b}w_{a}r_{b,1}r_{a,2}$
We have changed the product of sums into a sum of products. The number of terms grows exponentially with the number of periods. With every additional period, we double the number of terms in the two asset case.
The symbols $\{a,b\}$ can be combined into $2^{n}$ sequences of length $n$. Each such sequence corresponds to a term in the expanded $n$ period sum of products, indicating whether the return at time $t$ should be $a$’s or $b$’s return at that time.
If $n_{a}$ and $n_{b}$ indicate the number times each symbol occurs in a given sequence, the weight for the sequence is
(27) $w_{a}^{n_{a}}w_{b}^{n_{b}}$
The returns are more complex because they change over time. For now, think of them as an additional product that goes with each term.
What we now have is an exponential number of sequences that describe horse race style (mutually exclusive) investments on only one symbol per period. It’s straightforward to show that the weights $w_{a}^{n_{a}}w_{b}^{n_{b}}$ across these sequences add up to one. Effectively we have turned the multi-period compounding and rebalancing process into a single (non-rebalancing) weighted investment across an exponential number of multi-period investments. Imagine pre-allocating capital to an exponential number of processes and compounding serially within each siloed process.
Let $\mathcal{T}_{n}(P)$ be the set of sequences of length n with empirical frequencies given by
(28) $P$ $=\{p_{1},\dots,p_{m}\}$
(29) $\sum_{i=1}^{m}p_{i}$ $=1$
(30) $np_{i}$ $\in\mathbb{N}$
Each $\mathcal{T}_{n}(P)$ defines a type class of sequences. It is the set of all permutations of sequences of length $n$ with symbol frequencies matching $P$. This lets us write the total growth equation as a sum over all type classes
(31) $R_{n}(W)=\sum_{P}\prod_{i=1}^{m}w_{i}^{np_{i}}\sum_{s\in\mathcal{T}_{n}(P)}r_{s}$
where $r_{s}$ is the product of returns obtained by following sequence $s$, taking the return of whichever asset the sequence selects at each time period. This is the same exponential set of sequences as before, just grouped by type and with the common sequence weight $\prod_{i=1}^{m}w_{i}^{np_{i}}$ factored out per class.
As stated earlier, the weights sum to one across the full set of sequences. Each sequence weight represents a budget allocation to that particular multi-period compounding process that bets on one specific symbol per time period. A particular type class represents all permutations of a given sequence. Because all permutations in the class have the same weight, it’s not possible to wager more on a particular sequence than any other sequence in the set. The larger the set (higher the entropy), the less specific the wager. Cover used this fact to show that we can asymptotically achieve the same growth as an ideal adversary, one that knows the “rules” (the true odds) but not the future. The key insight is that even such an adversary is similarly restricted and can bet with high concentration only if the set is small (low entropy). By using the entropy of the sequence to size the wager, one can wager proportionally to the unknown optimal bet with only a polynomial penalty to make the wagers sum to one (the cost of universality). From a traditional finance perspective, this is a truly surprising result. The idea of optimal investing without any informational advantage seems paradoxical, but it’s really a demonstration of how powerful it is to apply information theory to the investing problem. Most people who are aware of Cover’s universal portfolio know it as a beautifully abstract and elegant yet likely impractical algorithm due to transaction costs. But as I’ll demonstrate, this isn’t just a scaffold for making universal portfolios. Cover’s idea to rewrite the traditional product of sums investing problem as a sum of products reveals the information structure of the general investing problem. It’s not just useful for Cover’s very abstract idea – it’s an essential tool that I argue makes the investing problem less opaque and more accessible to the techniques of information theory.
We will now demonstrate how the weight, which sums to one across all sequences, concentrates in the particular type class $P=W$, the set of all permutations of sequences with frequencies matching the portfolio weights. We first establish this using combinatorics and Stirling’s approximation. Because this approximation is only asymptotically valid, we then refine the argument using KL divergence to show how the total weight assigned to non-dominant sets declines exponentially with both divergence and time. Once this type class concentration is established we will use it to simplify the growth function to reveal the information theoretic terms in the investment process.
The size of $\mathcal{T}_{n}(W)$ is the number of permutations of sequences of length $n$ and with $n_{i}=w_{i}n$. The size of this set is given by the multinomial
(32) $\binom{n}{n_{1},n_{2},n_{3},\dots,n_{m}}=\frac{n!}{n_{1}!n_{2}!n_{3}!\dots n_{m}!}=\frac{n!}{\prod_{i=1}^{m}n_{i}!}$
For the $\mathcal{T}_{n}(W)$ type class, this is
(33) $\frac{n!}{\prod_{i=1}^{m}(nw_{i})!}$
Using Stirling’s Approximation
(34) $\log(n!)=n\log(n)-n+\mathcal{O}(\log(n))$
Ignoring the sublinear term (this approximation will be revisited below), we can rewrite the cardinality, using $\mathcal{T}_{n}(W)$ to denote both the cardinality and the set where it’s unambiguous:
(35) $\log(\mathcal{T}_{n}(W))$ $=\log(n!)-\log\left(\prod_{i=1}^{m}(nw_{i})!\right)$
(36) $=n\log(n)-n-\sum_{i=1}^{m}(nw_{i}\log(nw_{i})-nw_{i})$
(37) $=n\log(n)-n-\sum_{i=1}^{m}nw_{i}\log(n)-\sum_{i=1}^{m}nw_{i}\log(w_{i})+\sum_{i=1}^{m}nw_{i}$
(38) $=n\log(n)-n-n\log(n)\sum_{i=1}^{m}w_{i}-n\sum_{i=1}^{m}w_{i}\log(w_{i})+n\sum_{i=1}^{m}w_{i}$
(39) $=-n\sum_{i=1}^{m}w_{i}\log(w_{i})$
(40) $=n\sum_{i=1}^{m}w_{i}\log\left(\frac{1}{w_{i}}\right)$
(41) $=n\operatorname{H}(W)$
The cardinality is therefore
(42) $\mathcal{T}_{n}(W)=e^{n\operatorname{H}(W)}$
The weight assigned to a sequence $s$ is denoted by $\mathbf{w}_{s}$. This is the product of the portfolio weights for each period corresponding to the selected symbol. Because all sequences within the type class $\mathcal{T}_{n}(W)$ have the same symbol frequencies (they are permutations of each other), they share the same sequence weight. The sequence weight $\mathbf{w}_{s}$ for each sequence $s\in\mathcal{T}_{n}(W)$ is
(43) $\mathbf{w}_{s}=\prod_{i=1}^{m}w_{i}^{nw_{i}}$
Applying logs, this becomes
(44) $\log(\mathbf{w}_{s})$ $=\log\left(\prod_{i=1}^{m}w_{i}^{nw_{i}}\right)$
(45) $=\sum_{i=1}^{m}nw_{i}\log(w_{i})$
(46) $=-\sum_{i=1}^{m}nw_{i}\log\left(\frac{1}{w_{i}}\right)$
(47) $=-n\operatorname{H}(W)$
The weight for each sequence is therefore
(48) $\mathbf{w}_{s}=e^{-n\operatorname{H}(W)}$
At least for large n (with precise refinement later), the sum of this per-sequence weight across all of the sequences in the set is approximately
(49) $\sum_{s\in\mathcal{T}_{n}(W)}\mathbf{w}_{s}=e^{-n\operatorname{H}(W)}e^{n\operatorname{H}(W)}=1$
This indicates that the weight is concentrated in the set of sequences from the type class $P=W$. Again this is the set of all permutations of sequences with frequencies matching the portfolio weight vector. It is the uniform wager on all multi-period, horse-race style investments with frequencies matching the portfolio.
This concentration argument relied on Stirling’s Approximation. Although this is asymptotically accurate and it identifies the dominant type class, it overstates the finite log cardinality.
The intuition is that Shannon’s entropy, which characterizes an infinite stochastic source, is greater than the complexity of a finite object with the same frequency distribution. The finite case is formally in the domain of Kolmogorov Complexity Theory [10], but from a coding perspective, the reason is because knowing the full distribution and the number of symbols remaining reduces the uncertainty.
We can use KL divergence to more precisely bound the weight mass assigned to type classes that are different from $W$. The weight for any type class $P$ is $\prod_{i}w_{i}^{np_{i}}$ and the cardinality is approximately $|\mathcal{T}_{n}(P)|=e^{nH(P)}$. The log of the total density assigned to the set of sequences in the class is
(50) $\log\left(\prod_{i}w_{i}^{np_{i}}e^{nH(P)}\right)$ $=\sum(np_{i}\log\left(w_{i}\right))+n\operatorname{H}(P)$ $=-n\operatorname{H}(P,W)+n\operatorname{H}(p)$
(51) $=-n\operatorname{D_{KL}}(P\|W)$
The total weight concentrated on sequences in that set declines exponentially with the divergence from the portfolio weight vector $W$ and with the number of periods:
(52) $e^{-n\operatorname{D_{KL}}(P\|W)}$
This gives us confidence that we can ignore sequences that do not match the portfolio weights, with that confidence increasing exponentially with both the divergence (frequency mismatch) and the number of time periods. This lets us simplify the original growth equation, rewriting it as a single product rather than a complex sum over classes.
The original product expression
(53) $R_{n}(W)=\sum_{P}\prod_{i=1}^{m}w_{i}^{np_{i}}\sum_{s\in\mathcal{T}_{n}(P)}r_{s}$
can now be reduced to
(54) $e^{-n\operatorname{H}(W)}\sum_{s\in\mathcal{T}_{n}(W)}r_{s}$
Because $|\mathcal{T}_{n}(W)|=e^{n\operatorname{H}(W)}$, this is simply the average return across all multi-period sequence returns in the class of sequences that look like the portfolio (symbol frequencies equal to portfolio weights).
Now assume that $W^{*}$ is the optimal portfolio weight vector– the portfolio that maximizes the expected log return. As we showed above, almost all of its mass is assigned to the set of sequences with frequencies matching the portfolio weights. If we invest using any other weight vector $W\neq W^{*}$, the performance of this suboptimal portfolio is (ignoring all but the dominant set for the optimal portfolio)
(55) $R_{n}(W)$ $=\prod_{i=1}^{m}w_{i}^{nw_{i}^{*}}\sum_{s\in\mathcal{T}_{n}(W^{*})}r_{s}$
(56) $\log(R_{n}(W))$ $=\sum_{i=1}^{m}nw_{i}^{*}\log(w_{i})+\log\left(\sum_{s\in\mathcal{T}_{n}(W^{*})}r_{s}\right)$
(57) $=-n\operatorname{H}(W^{*},W)+\log\left(\sum_{s\in\mathcal{T}_{n}(W^{*})}r_{s}\right)$
(58) $=-nH(W^{*})-n\operatorname{D_{KL}}(W^{*}\|W)+\log\left(\sum_{s\in\mathcal{T}_{n}(W^{*})}r_{s}\right)$
[…]
The expected log return g(W) is therefore
[…]
The general case is more complicated than the classic Kelly or the horse race market problem. But we once again see that Kelly’s objective has factored the investing problem into the same three term decomposition: a money term (payoff / rules of the game), an entropy term (uncertainty) and a divergence term. As before, the first two terms don’t depend on our portfolio allocation. The only way to maximize the compounding growth rate is to minimize the drag or friction from the divergence which measures, in bits  [^1] Throughout, log denotes the natural logarithm. Results are stated in bits with the understanding that base conversion is a scaling constant., the difference between our distribution and the unknown true distribution. Therefore investing is, fundamentally, a compression problem.
5. Practical Applications
5.1. Backtesting
One of the challenges when backtesting and A/B testing of strategies is that market data is notoriously non-stationary. Even the log objective will be sensitive to extreme and atypical historical events. This makes the average difficult to interpret in a statistical sense when comparing strategies. But recall that the first two terms in equation 60 (money term and entropy term) do not depend on the allocation, so they are constant across strategies in a given backtest. The absolute value has no real statistical meaning, but the difference measures, in bits, the change in the divergence relative to some unknown ideal. This is reminiscent of Gibbs free energy in statistical mechanics.
[…]
(63) $g(W_{B})-g(W_{A})$ $=\operatorname{D_{KL}}(W^{*}\|W_{A})-\operatorname{D_{KL}}(W^{*}\|W_{B})$
5.2. MDL, Market Prediction, and Strategy Performance
Markets are high-dimensional, sparse, and non-stationary. This makes it very difficult to separate signal from noise. It is especially challenging for techniques based on statistical learning theory [20] which rely on resampling from a stationary stochastic source to detect overfitting and generalization. Although statistical learning theory is the dominant machine learning paradigm, it’s not the only one. Information theory offers alternative learning paradigms, including the MDL principle [8, 10]; see [16] for a discussion of prediction, compression, and the relationship between models and data. The key insight is that randomness is incompressible, so generalization is about finding a parsimonious, compressive description. The relationship between strategy performance and prediction performance is more direct when both are treated as information theoretic questions. In other words, it’s easier to draw a clear line between signal improvements and strategy improvements when both are measured in bits.
5.3. Winner Fraction and VC-style Investing
Kelly’s objective is convex. This means that there is a unique solution, and if we can model the joint return distribution, then we can solve the general case efficiently using convex optimization techniques.
But it’s often not feasible to accurately model the joint distribution. Consider the problem of venture capital style investing. The investor has to decide how much capital to allocate, often years before any real clarity on future cash flow.
We saw from equation 54 that the optimal portfolio concentrates on a particular type class. Finding that optimal type class is equivalent to solving the convex optimization problem which is not feasible without a precise joint probability distribution of future returns. Even without the full return distribution, the structure of this decomposition motivates a useful heuristic: winner fraction investing.
The principle is to invest in proportion to the fraction of universes where a given asset dominates the candidate set. In other words, the probability that a given startup “wins” relative to the rest of the contemplated startups. Although this is still nontrivial to predict, it is far less complicated than the full joint distribution – it does not require modeling the actual returns, just the probability of winning.
It will now be shown that this heuristic is approximately optimal. The lower the entropy, or the more extreme the “wins,” the better the approximation.
Let $W^{*}$ be the growth-maximum portfolio, and let $W^{\prime}$ be the vector of winner probabilities, in other words the probability that the $i$th startup dominates the candidate set.
(64) $\frac{R_{n}(W^{*})}{R_{n}(W^{\prime})}=\frac{e^{-n\operatorname{H}(W^{*})}\sum_{s\in\mathcal{T}_{n}(W^{*})}r_{s}}{e^{-n\operatorname{H}(W^{\prime})}\sum_{s\in\mathcal{T}_{n}(W^{\prime})}r_{s}}$
The numerator and denominator represent averages over their respective type classes.
Recall that each sequence is an exclusive bet on only one asset per time period. A type class represents all sequence permutations with a given frequency matching the portfolio weight vector. The largest return in the winner fraction set is constructed by stringing together the dominant asset per period. This sequence must be in the winner fraction set. Let $r_{\text{max}}$ be the return for that sequence. By construction, it must also be at least as large as every sequence in the optimal set $\mathcal{T}_{n}(W^{*})$. Therefore
(65) $\forall_{s\in\mathcal{T}_{n}(W^{*})}\quad\frac{r_{s}}{r_{\text{max}}}\leq 1$
(66) $\frac{R_{n}(W^{*})}{R_{n}(W^{\prime})}$ $\leq\frac{e^{-n\operatorname{H}(W^{*})}\sum_{s\in\mathcal{T}_{n}(W^{*})}r_{s}}{e^{-n\operatorname{H}(W^{\prime})}r_{\text{max}}}$
(67) $\leq\frac{\frac{1}{|\mathcal{T}_{n}(W^{*})|}\sum_{s\in\mathcal{T}_{n}(W^{*})}\frac{r_{s}}{r_{\text{max}}}}{e^{-nH(W^{\prime})}}$
(68) $\leq e^{nH(W^{\prime})}$
This shows the ratio of the optimal to the heuristic solution.
Taking limits to find expectation, we see
(69) $g(W^{*})-g(W^{\prime})\leq\lim_{n\to\infty}\frac{\log(e^{nH(W^{\prime})})}{n}=H(W^{\prime})$
The difference between growth-optimal portfolio and the winner-fraction portfolio is bounded by the entropy of the winner-fraction distribution. That entropy is itself bounded by the cardinality of the candidate set, so this heuristic is especially useful when the candidate set is small or when the return distribution is extreme. To my knowledge, this heuristic and proof of the entropy bound are both original.
6. Conclusion
The Kelly criterion describes how to allocate capital in simple games of chance with known odds and edge. It is a closed form optimal solution to Kelly’s log wealth objective. But it turns out that Kelly’s objective is itself optimal among all objectives. And this fact reframes the investing problem itself.
In finance, we often think of risk and reward as two different axes: for a given amount of risk, we maximize return. This goes back to the 1940’s and Markowitz’s Modern Portfolio Theory (MPT) [12]. A concern with MPT is that markets are heavy-tail and so diversification (averaging) may not reduce risk (variance) [17]. This is fundamentally a statistical concern. But even if markets were Gaussian, it would still be suboptimal because Kelly’s objective is uniquely optimal. To justify that, we have to turn from statistics to information theory.
Instead of thinking of risk and reward as two different axes, Kelly proposed maximizing a single objective, the logarithm of wealth over uncertainty. In other words the expectation or the probability weighted average log return. Samuelson argued against this idea, successfully keeping it out of mainstream economics.
We now know because of Tom Cover’s work at Stanford that this was not a subjective, arbitrary utility function as Samuelson believed. It is the objectively optimal way to invest. It is competitively optimal in a game-theoretic sense, assuming neither party has an information advantage. It dominates over almost every market sequence. Even over the short term.
With the help of a transformation from Cover’s universal portfolio algorithm, we now understand what is so powerful about Kelly’s objective. It reveals the information structure of the investing problem, factoring it into three terms: a money term (payoff / rules of the game), an entropy term (uncertainty) and a divergence term. The only way to maximize compounding growth rate is to minimize the friction from that divergence term. Fundamentally, investing is a compression problem.